OpenStudy (anonymous):

if a quadratic equation (example) is like x^2 + 3x - (k-3), in the quadratic formula, would the discriminant be 3^2 - 4(1)(k-3) or 3^2 - 4(1)(-k+3)?

OpenStudy (anonymous):

disc= 3^2 - 4(1)[-(k-3)]

OpenStudy (anonymous):

= 9+4(k-3)

OpenStudy (anonymous):

=4k-3

OpenStudy (amistre64):

the 2nd one...

OpenStudy (anonymous):

thank so much, both of you

OpenStudy (amistre64):

-(k-3) = c

OpenStudy (anonymous):

best of luck with ur maths...

OpenStudy (anonymous):

thanks, i'll be back with more questions later haha

OpenStudy (anonymous):

sure ...

OpenStudy (anonymous):

fan me if u could..cheers

OpenStudy (anonymous):

i already am :)

OpenStudy (amistre64):

unfan me if you want ;) your choice lol